The integral curves of dx/a = dy/b = dz is


1
Expert's answer
2021-09-21T18:14:29-0400

Let us find the integral curves of "\frac{dx}{a} = \frac{dy}{b} = dz." It is equivalent to "dz=\frac{dx}{a}" and "dz= \frac{dy}{b}." Then"\int dz=\int\frac{dx}{a}" and "\int dz= \int\frac{dy}{b}." It follows that "z=\frac{x}{a} +C_1" and "z= \frac{y}{b}+C_2." Therefore, "z=\frac{x}{a} +C_1=\frac{y}{b}+C_2" is the required integral curves.